OpenStudy (anonymous):

I got an answer but it seems that its wrong please help! limit horizental question

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} 5x+10/x-11 *6x-10/-x-4\]

OpenStudy (anonymous):

A function is said to have a horizontal asymptote if either the limit at infinity exists or the limit at negative infinity exists. Show that each of the following functions has a horizontal asymptote by calculating the given limit.

OpenStudy (anonymous):

I got 30!

OpenStudy (noelgreco):

Some parentheses would really help.

OpenStudy (anonymous):

mmm.._?

OpenStudy (anonymous):

i agree I can't tell if everything is being divided and multiplied or just certain things

OpenStudy (anonymous):

ok, you want to see how i got this answer?

OpenStudy (anonymous):

The answer is -30 Apply partial fraction decomposition to each fraction on the LHS. http://www.purplemath.com/modules/partfrac.htm \[\frac{5x+10}{x-11}*\frac{6x-10}{-x-4}=\left(\frac{65}{x-11}+5\right) \left(\frac{34}{x+4}-6\right) \]

OpenStudy (anonymous):

Take the derivative of each numerator and each denominator and evaluate the result. \[\frac{D[5x+10,x]}{D[x-11,x]}*\frac{D[6x-10,x]}{D[-x-4,x]}=\frac{5}{1}*\frac{6}{(-1)}=-30 \]D[ f[x], x ] is Mathematica notation for " the derivative of f(x) with respect to x"